if the roots of the equation x^2+3x+2=0 and x^2-x+lambda =0 are in the same ratio then the value of lambda is

Grade:11

1 Answers

Shailendra Kumar Sharma
6 years ago
Roots of X2+3x+2 be X=-1 and X= -2 So ratio will be 1/2 Assume roota of x2-x+ lambda be a and 2a Then a+2a = -(-1/1)= 1So 3a= 1 So a=1/3And products of roots a*2a = lambda/1So lambda =2a^2 =2/9